Electromagnetic radiation of wavelength 96.4 nm was projected on a zinc plate with a 6.01×10-19 J of work function. Calculate the kinetic energy of an emitted electron.

a. -6.01×10-19 J

b. 1.46×10-18 J

c. 28.9 J

d. 2.66×10-18 J

e. 4.69×1048 J
Answer #1

    = 96.4nm

       = 96.4*10^-9m

W   = 6.01*10^-19 J

KE   = hc/ - W

        = 6.625*10^-34 * 3*10^8/(96.4*10^-9) - 6.01*10^-19

       = 2.06*10^-18 - 6.01*10^-19

       = 2.06*10^-18 - 0.601*10^-18

       = 1.46*10^-18 J

b. 1.46×10-18 J >>>>answer